# arXiv:2403.09539v2[cs.CL] 15 Mar 2024

### v2

@@ 2 LLM outputs are restricted to a low-dimensional linear space
The rest of this paper discusses the implications and applications of this mathematical fact for API-protected LLMs, starting with methods for finding the LLM image given a restrictive API, then using the LLM image for various purposes.
3 Fast, full outputs from API-protected LLMs
There are several uses for full-vocab LLM outputs, however most LLM APIs do not return full outputs. This is likely because full outputs are large and expensive to send over an API, but could also be to prevent abuse of APIs, such as using these outputs to distill models (Hinton et al., 2015; Mukherjee and Awadallah, 2019; Hsieh et al., 2023) or discover proprietary information (see Section 4). In their paper, Morris et al. (2023) give an algorithm for recovering these full outputs from restricted APIs by taking advantage of a common API option that allows users to add bias to the logits for specific tokens. The algorithm they describe requires $O(v\log\frac{1}{\epsilon})$ calls to the API top obtain one full output with precision $\epsilon$.
We give an improvement that theoretically obtains full outputs in $v/k$ API calls for APIs that return the log-probability of the top-$k$ tokens. We find that this improved algorithm suffers from numerical instability, and give a numerially stable algorithm that obtains full outputs in v/(k1)v/(k-1) API calls.
Next, we give a practical algorithm for dealing with stochastic APIs that randomly choose outputs from a set of $n$ possible outputs. This algorithm allows the collection of full outputs in less than nv/(k2)nv/(k-2) API calls on average.
Finally, we reduce the number of API calls in all of the above algorithms from $O(v)$ to $O(d)$ by adding in a preprocessing step to find the low-dimensional image of the LLM. This speedup makes obtaining full LLM outputs up to 100 times faster and cheaper, depending on the LLM. Table 2 gives an overview of our algorithms with back-of-the envelope cost estimates for a specific LLM.
Table 2: A summary of our proposed algorithms, with estimates for the number of API calls required per output, and the price of acquiring the model image. Estimates are based on a gpt-3.5-turbo-like API LLM with $v=100\text{K}$, $d=4096$, $\epsilon=10^{-6}$, $k=5$, $b_{\max}=100$, and $n=4$. Note that the $O(d)$ algorithm cannot be used to obtain the LLM image, since it relies on having LLM image as a preprocessing step.
Algorithm Complexity API calls per output Image price
Logprob-free $O(v\log\frac{1}{\epsilon})$ 800K \$16,384
With logprobs $v/k$ 20K \$410
Numerically stable v/(k1)v/(k-1) 25K \$512
Stochastically robust nv/(k2)nv/(k-2) 133K \$2,724
LLM Image $O(d)$ 1K–32K -
3.1 Fast, full outputs from APIs with logprobs
Our goal is to recover a full-vocabulary next-token distribution $\boldsymbol{p}\in\Delta_{v}$ from an API-protected LLM. We will assume that the API accepts a prompt on which to condition the distribution, as well as a list of up to $k$ tokens and a bias term $b\leq b_{\max}$ to add to the logits of the listed tokens before applying the softmax function. The API returns a record with the $k$ most likely tokens and their probabilities from the biased distribution. For instance, querying the API with $k$ maximally biased tokens, which without loss of generality we will identify as tokens $1,2,\ldots,k$, yields the top-$k$ most probable tokens from the biased distribution $\boldsymbol{p}^{\prime}=\operatorname{softmax}(\boldsymbol{\ell}^{\prime})$ where
$\ell^{\prime}_{i}=\begin{cases}\ell_{i}+b_{\max}&i\in\{1,2,\ldots,k\}\\ \ell_{i}&\text{otherwise}\end{cases}$
Assuming that the logit difference between any two tokens is never greater than $b_{\max}$, these top-$k$ biased probabilities will be $p^{\prime}_{1},p^{\prime}_{2},\ldots,p^{\prime}_{k}$. For each of these biased probabilites $p^{\prime}_{i}$, we can solve for the unbiased probability as
$p_{i}=\frac{p^{\prime}_{i}}{\exp b_{\max}-\exp b_{\max}\sum_{j=1}^{k}p^{\prime}_{j}+\sum_{j=1}^{k}p^{\prime}_{j}}$ (2)
(proof in the Appendix A.1). Thus, for each API call, we can obtain the unbiased probability of $k$ tokens, and obtain the full distribution in $v/k$ API calls.
3.2 Numerically stable full outputs from APIs
In practice, the algorithm described in Section 3.1 suffers from severe numerical instability, which can be attributed to the fast-growing exponential term $\exp b_{\max}$, and the term $\sum_{j=1}^{k}p^{\prime}_{j}$ which quickly approaches 1. We can eliminate the instability by sacrificing some speed and using a different strategy to solve for the unbiased probabilities. Without loss of generality, let $p_{v}$ be the maximum unbiased token probability. This can be obtained by querying the API once with no bias. If we then query the API and apply maximum bias to only tokens $1,2,\ldots,1-k$, then the API will yield $p^{\prime}_{1},p^{\prime}_{2},\ldots,p^{\prime}_{k-1}$ and $p^{\prime}_{v}$. We can then solve for the unbiased probabilities of the k1k-1 tokens
$p_{i}=\exp(\log p^{\prime}_{i}-b_{\max}-\log p^{\prime}_{v}+\log p_{v})$ (3)
(proof in Appendix A.2). By finding k1k-1 unbiased token probabilities with every API call, we obtain the full output in v/(k1)v/(k-1) calls total.
3.3 Full outputs from stochastic APIs
Each of the above algorithms assume that the API is deterministic, i.e., the same query will always return the same output. This is not always the case. For instance, we find that OpenAI’s LLM APIs are stochastic. While this would seem to doom any attempt at obtaining full outputs from the LLM, we find that certain types of stochasticity can be dealt with. In particular, we model OpenAI’s stochastic behavior as a collection of $n$ outputs $\boldsymbol{p}^{(1)},\boldsymbol{p}^{(2)},\ldots,\boldsymbol{p}^{(n)}$ from which the API randomly returns from. This might be the result of multiple instances of the LLM being run on different hardware which results in slightly different outputs. Whichever instance the API returns from determines which of the $n$ outputs we get. In order to determine which of the outputs the API returned from, assume without loss of generality that $p^{(i)}_{v-1}$ is the second highest token probability for all $i$, and observe that
$\log p^{(i)}_{v}-\log p^{(i)}_{v-1}=\log p^{\prime(i)}_{v}-\log p^{\prime(i)}_{v-1}$ (4)
for all outputs $\boldsymbol{p}^{(i)}$ and unbiased outputs $\boldsymbol{p}^{\prime(i)}$ where tokens $v$ and v1v-1 are not biased (proof in Appendix A.3). Therefore, by biasing only k2k-2 tokens for each call, the API will return $p^{\prime(i)}_{1},p^{\prime(i)}_{2},\ldots,p^{\prime(i)}_{k-2}$ as well as $p^{\prime(i)}_{v}$ and $p^{\prime(i)}_{v-1}$ for some $i$, and we can determine which output $\boldsymbol{p}^{(i)}$ the result comes from by using $\log p^{\prime(i)}_{v}-\log p^{\prime(i)}_{v-1}$ as an identifier. Thus, after an average of nv/(k2)nv/(k-2) calls to the API we can collect the full set of probabilities for one of the outputs.
3.4 Ultra-fast full outputs using the LLM image
So far, the dominating factor in each algorithm’s runtime is $v$. We now introduce a preprocessing step that takes advantage of the low-dimensional LLM output space to obtain $O(d)$ versions of all the above algorithms. Since $d\ll v$ for many modern language models, this modification can result in multiple orders of magnitude speedups, depending on the LLM. For instance, the speedup for an LLM like pythia-70m would be $100\times$. The key to this algorithm is the observation from Section 2 that $d$ linearly independent outputs from the API constitute a basis for the whole output space (since the LLM’s image has dimension $d$). We can therefore collect these outputs
$\boldsymbol{P}=\begin{bmatrix}\boldsymbol{p}^{1}&\boldsymbol{p}^{2}&\cdots&\boldsymbol{p}^{d}\end{bmatrix}\in\Delta_{v}^{d}$
as a preprocessing step in $O(vd)$ API calls using any of the above algorithms and $d$ unique prompts, and then use these to reconstruct the full LLM output after only $O(d)$ queries for each subsequent output.
To get a new full output $\boldsymbol{p}$, use any of the above algorithms to obtain $p_{1},p_{2},\ldots,p_{d}$. Next, we will use the additive log ratio ($\operatorname{alr}$) transform, which an isomorphism $\Delta_{v}\to\mathbb{R}^{v-1}$ and is defined as
$\operatorname{alr}(\boldsymbol{p})=\left(\log\frac{p_{2}}{p_{1}},\log\frac{p_{3}}{p_{1}},\ldots,\log\frac{p_{v}}{p_{1}}\right)$
$\operatorname{alr}(\boldsymbol{P})=\begin{bmatrix}\operatorname{alr}(\boldsymbol{p}^{1})&\operatorname{alr}(\boldsymbol{p}^{2})&\cdots&\operatorname{alr}(\boldsymbol{p}^{1})\end{bmatrix}\in\mathbb{R}^{(v-1)\times d}$
$\operatorname{alr}^{-1}(\boldsymbol{x})=\frac{1}{1+\sum_{i=1}^{v-1}\exp x_{i}}\cdot\left(1,\exp x_{1},\exp x_{2},\ldots,\exp x_{v-1}\right).$
Thus we can retrieve $\boldsymbol{p}$ in only $O(d)$ API queries.
This $(v/d)\times$ speedup makes any method that relies on full model outputs significantly cheaper. This includes model stealing (Tramèr et al., 2016) which attempts to learn a model that exactly replicates the behavior of a target model.
4 Discovering the embedding size of API-protected LLMs
Without knowing the exact architectural details, other than assuming the generic output layer described in Figure 2, it is possible to infer the embedding size $d$ of an API-protected language model from its outputs alone. This is a direct result of the fact that the model outputs occupy a $d$-dimensional subspace of $\Delta_{v}$, and therefore collecting $d$ linearly independent outputs $\boldsymbol{p}_{1},\boldsymbol{p}_{2},\ldots,\boldsymbol{p}_{d}\in\Delta_{v}$ from the LLM will form a basis for the LLM’s image, meaning that all subsequent model outputs $\boldsymbol{p}_{d+1}$ will be a linear combination of the first $d$ outputs. Assuming that model outputs are full rank, i.e., all collections of $d$ prompts are linearly independent, and all collections of $d+1$ prompts are linearly dependent, we can discover the value of $d$ by collecting outputs one at a time until the number of linearly independent outputs in the collection stops increasing, which will occur when we have collected $d+1$ prompts. In practice, we find that this full-rank assumption tends to hold well enough when over-collecting outputs, e.g., collecting $d+1000$ outputs.
To find the dimension of a space spanned by model outputs, we use the fact that a matrix with $d$ linearly independent columns will have $d$ non-zero singular values. We can therefore plot the singular values of the matrix
$\boldsymbol{L}=\operatorname{clr}(\boldsymbol{P})=\begin{bmatrix}\operatorname{clr}(\boldsymbol{p}_{1})&\operatorname{clr}(\boldsymbol{p}_{2})&\cdots&\operatorname{clr}(\boldsymbol{p}_{n})\end{bmatrix}$
and observe the index at which the magnitude of the singular values drops to zero, which will occur at index dd. [^2] We will loosely refer to $\boldsymbol{L}$ as the logit matrix, though these are technically the logits projected onto the nearest point on $U_{v}$. In practice, due to precision issues, the magnitudes may not drop all the way to zero.
Figure 4: The singular values of outputs from LLMs with various known and unknown embedding sizes $d$. For each model with known embedding size, there is a clear drop in magnitude at singular value index $d$, indicating the embedding size of the model. Using this observation, we can guess the embedding size of gpt-3.5-turbo.
To validate our method, we collect next-token distributions (conditioned on unique, 1-token prompts) from several open-source LLMs from the Pythia family (Biderman et al., 2023) with embedding sizes ranging from 512 to 1024. For all these models, we find that the singular values of the resulting output matrix drop dramatically precisely at the index corresponding to the embedding size of the model, as shown in the first three plots of Figure 2.
To demonstrate our method’s effectiveness against API-protected LLMs, we use our stochastically robust nv/(k2)nv/(k-2) algorithm to collect nearly 6,000 next-token distribution outputs from gpt-3.5-turbo [^3] We specifically use model version 0125, accessed February 1–19, 2024., a popular API-protected LLM whose embedding size is not publicly disclosed. We find that the singular values for these outputs drop dramatically between index 4,600 and 4,650, indicating that the embedding size of this model is at most this size. This predicted embedding size is somewhat abnormal, since LLM embedding sizes are traditionally set to powers of two (or sums of powers of two). If this were the case for gpt-3.5-turbo, it would be reasonable to guess that the embedding size is $2^{12}=4096$ or $2^{12}+2^{9}=4608$, the former of which we think is most likely. We predict that our raw estimate of 4,600-4,650 is an overestimate of the true embedding size, since any abnormal outputs due to errors (whether in our own code or OpenAI’s) would only increase the dimensionality of the observed output space. In particular, if we inadvertently collected 504 corrupted outputs, then a model with embedding size 4,096 would appear to have an embedding size of 4,600.
Extrapolating further, assuming that gpt-3.5-turbo has a similar architecture to most transformer-based LLMs, it is likely that the number of parameters in gpt-3.5-turbo is around 7 billion. This estimation is based on the fact that most known transformer-based LLMs with embedding size 4,096 have $\approx 7$ billion parameters. Any other parameter count would result in either abnormally narrow or wide neural networks which are not known to have favorable performance. An important exception to this rule-of-thumb, however, would be the increasingly popular “mixture-of-experts” architecture which tends to have many more parameters per embedding dimension.
Previous estimates of gpt-3.5-turbo’s parameter count based on hearsay have generally exceeded 7B (Singh et al., 2023), however given the periodically updating versions and decreasing cost of inference with this model, it is possible that its size and architecture has changed over time. Fortunately, our method can be used to monitor these updates over time, alerting end-users when LLM providers make architectural changes to their model, specifically updates to the embedding size and/or parameter count.
5 Identifying LLMs from a single output
[…]
The image of two different LLMs, even different checkpoints from the same LLM, are largely disjoint. As shown in Figure 5, the logit output from a late-training checkpoint of Pythia-70M lies uniquely in the image of the checkpoint, and not in the image of the preceding or following checkpoints. We also verify that the output does not lie in the image of any checkpoints from a similar LLM trained on deduplicated data, nor any checkpoint from a larger LLM trained on the same data. This suggests that the image of a LLM is highly sensitive to small parameter changes. Intuitively this makes sense, since the output of one model will only be in the image of another model in the extremely unlikely event that it happens to lie on the low-dimensional (<d<d) intersection of the models’ images. Mathematically the dimension of the models’ images’ intersection is small since the intersection $U\cap V$ of two subspaces $U$ and $V$ that are not subsets of one another has dimension $\dim U+\dim V-\dim(U\cup V)$, which implies that $\dim(U\cap V)<\min\{\dim U,\dim V\}$ (since $\dim(U\cup V)>\max\{\dim U,\dim V\}$). Thus, it is possible to determine precisely which LLM produced a particular output, using only API access to a set of LLMs and without knowing the exact inputs to the model. In this way, the model’s image acts as a signature for the model, i.e., a unique model identifier.
This finding has implications for LLM provider accountability. For instance, if provider $A$ suspects provider $B$ of stealing their proprietary model, provider $A$ could check provider $B$’s model signature to see if it matches their own. It would be extremely unlikely that the signatures match if the LLMs are not the same. Similarly, if a provider claimed to be serving a proprietary model while in actuality attempting to profit off of an open-source model with a non-commercial license, an auditor again could discover this by checking the LLM signature. This test is somewhat one-sided, however, since a clever provider may do a small amount of additional fine-tuning to change their stolen LLM’s image.
5.1 Detecting and disambiguating LLM updates
@@
Another practical application of highly sensitive LLM images is to detect and distinguish minor and major model updates to API-protected LLMs. Using the above test, we can tell whether or not the LLM image remains the same, even if the logit outputs change. This phenomenon would correspond to a partial model update, where some part of the model changes, but the softmax matrix remains the same. An example of this would be if the LLM has a hidden prefix added to all prompts and this hidden prefix changes: the model’s outputs will change but the image will remain the same. Table 3 gives an overview for how to interpret various combinations of detectable API changes, This information is useful for monitoring unannounced changes in API LLMs.
Table 3: Implications of image/logit changes.
@@
Logit change, no image change Hidden prompt change or partial finetune with frozen output layer
Low-rank image change (See § 5.2) LoRA update
@@ 5.2 Detecting LoRA updates
We speculate that it is possible to gain even higher granularity information on LLM updates of certain types, such as LoRA (Hu et al., 2022) updates. LoRA is a popular parameter-efficient fine-tuning method which adjusts model weights with a low-rank update $\boldsymbol{A}\boldsymbol{B}$ where $\boldsymbol{A}\in\mathbb{R}^{v\times r}$ and $\boldsymbol{B}\in\mathbb{R}^{r\times d}$ so that the softmax matrix $\boldsymbol{W}\in\mathbb{R}^{v\times d}$ becomes $\boldsymbol{W}+\boldsymbol{A}\boldsymbol{B}$. We speculate that it is possible to detect these types of updates by collecting LLM outputs before ($\boldsymbol{L}\in\mathbb{R}^{v\times d}$) and after ($\boldsymbol{L}^{\prime}\in\mathbb{R}^{v\times d}$) the update and decomposing them as $\boldsymbol{W}\boldsymbol{H}=\boldsymbol{L}$ and $(\boldsymbol{W}+\boldsymbol{A}\boldsymbol{B})\boldsymbol{H}=\boldsymbol{L}^{\prime}$ where $\boldsymbol{H},\boldsymbol{H}^{\prime}\in\mathbb{R}^{d\times d}$ such that $\operatorname{im}(\boldsymbol{W}\boldsymbol{H})=\operatorname{im}(\boldsymbol{W}+\boldsymbol{A}\boldsymbol{B})$. If such a decomposition is found, then it appears likely that the weights received a low-rank update of rank $r$. We leave it to future work to determine whether finding such a decomposition is sufficient to conclude that a LoRA update took place, as well as to find an efficient algorithm for finding such a decomposition.
6 More applications
Access to the LLM’s image can lead to many other capabilities, some of which we discuss below. We leave further investigation of these methods for future work.
@@ 6.1 Finding unargmaxable tokens
Due to the low-rank constraints on LLM outputs, it is possible that some tokens become unargmaxable (Demeter et al., 2020; Grivas et al., 2022), i.e., there is a token $i$ such that the constraints disallow any output $\boldsymbol{p}$ with $\arg\max_{i}p_{i}=i$. This happens when the LLM’s embedding representation of $i$ lies within the convex hull of the other tokens’ embeddings. Previously, finding unargmaxable tokens appeared to require full access to the softmax matrix $\boldsymbol{W}$. Interestingly, we find that it is possible to identify unargmaxable tokens using only the LLM’s image, which our method is able to recover. This technique allows API customers to find quirks in LLM behavior such as tokens that the model is unable to output (at least under greedy decoding).
@@ 6.2 Recovering the softmax matrix from outputs
Since our method gives us access to the image of the output layer, and not the output layer parameters, we investigate the how one might reconstruct the output later parameters, either exactly or approximately. We hypothesize that LLM embeddings generally lie near the surface of a hypersphere in $\mathbb{R}^{d}$ with a small radius $r$. We see evidence of this in the fact that the Pythia LLM embedding norms are all small and roughly normally distributed, as shown in Figure 6.
Assuming this holds for any LLM, we can attempt to recover $\boldsymbol{W}$ up to a rotation by simplifying the assumption into a constraint that all embeddings must have a magnitude of 1. Then, given a matrix $\boldsymbol{L}\in\mathbb{R}^{n\times v}$ of model logits, we can find $\boldsymbol{W}$ (up to a rotation) by finding a decomposition $\boldsymbol{W}\boldsymbol{H}=\boldsymbol{L}$ such that for all $i$, $\lVert W_{i}\rVert_{2}=r$. This solution may also be approximated by finding the singular value decomposition $\boldsymbol{W}\boldsymbol{\Sigma}\boldsymbol{V}^{\top}$ of $\boldsymbol{L}$, though it is likely that all rows of this $\boldsymbol{W}$ will have magnitude less than $1$ and they are not guaranteed to be normally distributed.
6.3 Improved LLM inversion
Morris et al.’s (2023) recent approach to recovering hidden prompts (i.e., an additional prefix added to LLM input, not shown to the user) uses full LLM API outputs to train an “inversion” model that generates the prefix conditioned on the LLM’s full logprob vector $\log\boldsymbol{p}$. In addition to our algorithm from Section 3.4 making this much cheaper to implement, we also propose a methodological change to their procedure. In particular, to deal with the size mismatch between the LLM’s vocabulary size $v$ and the inversion model’s embedding size $d_{\mathrm{inv}}$, Morris et al. (2023) pad and reshape the vector into a $\lceil v/d_{\mathrm{inv}}\rceil$-length sequence of embeddings in $\mathbb{R}^{d_{\mathrm{inv}}}$. This transformation is somewhat unnatural, and requires that the inversion model only be conditioned on a single output.
We propose instead to take advantage of our knowledge of the LLM’s image to obtain a linearly lossless representation of the LLM output in $\mathbb{R}^{d}$, which is much closer to the inversion model size, then use an MLP to project this representation onto $\mathbb{R}^{d_{\mathrm{inv}}}$ to feed into the inversion model. Formally, after obtaining $\boldsymbol{P}$ and $\boldsymbol{p}$ from the API, we use the unique solution $\boldsymbol{x}$ to $\operatorname{clr}(\boldsymbol{P})\boldsymbol{x}=\operatorname{clr}(\boldsymbol{p})$ as the input to the inversion model. This modification has an additional advantage: instead of conditioning on a single output $\boldsymbol{p}$, the LLM can be used to generate a short sequence of outputs $\boldsymbol{p}^{1},\boldsymbol{p}^{2},\ldots,\boldsymbol{p}^{n}$ which can be fed into the inversion model as $\boldsymbol{x}^{1},\boldsymbol{x}^{2},\ldots,\boldsymbol{x}^{n}$ where $\operatorname{clr}(\boldsymbol{P})\boldsymbol{x}^{i}=\operatorname{clr}(\boldsymbol{p}^{i})$. We leave the implementation and evaluation of this proposed modification for future work. This technique can be applied generally to any method that takes LLM outputs as input.

### v3

@@ 6 Attributing model outputs and auditing model updates
This finding has implications for LLM auditing (Mökander et al., 2023), such as obtaining granular information about updates to API-protected LLMs. For instance, if the LLM image remains the same but the logit outputs change, this would indicate a partial model update where some part of the model changes but the softmax matrix remains the same. This might happen when the LLM has a hidden prefix added to all prompts and this hidden prefix changes, or when some part of the model was updated while leaving the softmax matrix unchanged. table˜2 gives an overview for how to interpret various combinations of detectable API changes.
@@
Logit change, no image change Hidden prompt change or partial finetune
Low-rank image change (§7.1) LoRA update
@@
Table 2: Implications of image/logit changes.
Using the same principle, one can also detect malicious use of open-source LLMs. If a provider attempts to profit off of an open-source model with a non-commercial license, an auditor can reveal the scheme by checking the LLM image, even if the provider used a hidden prompt to mask the model identity. It would be unlikely for images of two different LLMs to match, unless purposely designed to do so. Note that this test is one-sided, and malicious providers may slightly fine-tune an LLM to avoid detection.
7 More Applications
Access to the LLM’s image unlocks several additional capabilites, some of which we review below. We leave further investigation of these methods for future work.
@@ 7.1 Detecting LoRA updates
Access to the LLM image can afford even finer granularity insight into LLM updates. For instance, as LoRA (Hu et al., 2022) is a popular parameter-efficient fine-tuning method which adjusts model weights with a low-rank update $\boldsymbol{A}\boldsymbol{B}$ where $\boldsymbol{A}\in\mathbb{R}^{v\times r}$ and $\boldsymbol{B}\in\mathbb{R}^{r\times d}$ so that the softmax matrix $\boldsymbol{W}\in\mathbb{R}^{v\times d}$ becomes $\boldsymbol{W}+\boldsymbol{A}\boldsymbol{B}$. It may be possible to detect these types of updates by collecting LLM outputs before ($\boldsymbol{L}\in\mathbb{R}^{v\times d}$) and after ($\boldsymbol{L}^{\prime}\in\mathbb{R}^{v\times d}$) the update and decomposing them as $\boldsymbol{W}\boldsymbol{H}=\boldsymbol{L}$ and $(\boldsymbol{W}+\boldsymbol{A}\boldsymbol{B})\boldsymbol{H}^{\prime}=\boldsymbol{L}^{\prime}$ where $\boldsymbol{H},\boldsymbol{H}^{\prime}\in\mathbb{R}^{d\times d}$. If such a decomposition is found then it is likely that the weights received a low-rank update. We leave it to future work to find an efficient algorithm for this decomposition.
@@ 7.2 Finding unargmaxable tokens
Due to the low-rank constraints on LLM outputs, it is possible that some tokens become unargmaxable (Demeter et al., 2020; Grivas et al., 2022), i.e., the token always has less probability than some other token. This happens when the LLM’s embedding representation of the token lies within the convex hull of the other tokens’ embeddings. Previously, finding unargmaxable tokens appeared to require full access to the softmax matrix $\boldsymbol{W}$, however, it is possible to identify unargmaxable tokens using only the LLM’s image, which our method is able to recover. This technique allows API clients to find tokens that the model is unable to output and thereby elicit unexpected model behavior.
@@ 7.3 Recovering the softmax matrix from outputs
One might use the image to approximately reconstruct the output layer parameters. We hypothesize that LLM embeddings generally lie near the surface of a hypersphere in $\mathbb{R}^{d}$ with a small radius $r$. We see evidence of this in the fact that the Pythia LLM embedding norms are all small and roughly normally distributed, as shown in figure˜6 in the Appendix. We can attempt to recover $\boldsymbol{W}$ up to a rotation by assuming that all embeddings must have unit magnitude, then, given a matrix $\boldsymbol{L}\in\mathbb{R}^{n\times v}$ of model logits, we can find $\boldsymbol{W}$ (up to a rotation) by finding a decomposition $\boldsymbol{W}\boldsymbol{H}=\boldsymbol{L}$ such that for all $i$, $\lVert W_{i}\rVert_{2}=1$. This solution may also be approximated by finding the singular value decomposition $\boldsymbol{W}\boldsymbol{\Sigma}\boldsymbol{V}^{\top}$ of $\boldsymbol{L}$, though rows of this $\boldsymbol{W}$ will have magnitude less than $1$.